# Fast training and sampling of Restricted Boltzmann Machines

### v2

@@ Appendix H Overfitting and privacy loss as quality indicators
In this section, we examine the quality of the samples generated, regarding overfitting and privacy criteria which have been defined for genomic data in particular. We look at this on the models trained with PCD with and without pre-training. We focus on the human genome dataset, as shown in Fig. 1–C, to evaluate the ability of various state-of-the-art generative models to generate realistic fake genomes while minimizing privacy concerns (i.e., reducing overfitting). Recent studies (Yelmen et al., 2021; 2023) have thoroughly investigated this for a variety of generative models. Both studies concluded that the RBM was the most accurate method for generating high-quality and private synthetic genomes. The comparison between models relies primarily on the Nearest Neighbor Adversarial Accuracy ($AA_{\rm TS}$) and privacy loss indicators, introduced in Ref. Yale et al. (2020), which quantify the similarity and the level of ”privacy” of the data generated by a model w.r.t. the training set. We have $AA_{\rm TS}=\frac{1}{2}\bigl{(}AA_{\rm True}+AA_{\rm Synth}\bigr{)}$ where $AA_{\rm True}$ [resp. $AA_{\rm Synth}$] are two quantities in $[0,1]$ obtained by merging two sets of real and synthetic data of equal size $N_{s}$ and measuring respectively the frequency that a real [rep. synthetic] has a synthetic [resp. real] as nearest neighbor. If the generated samples are statistically indistinguishable from real samples, both frequencies $AA_{\rm True}$ and $AA_{\rm Synth}$ should converge to 0.5 at large ${N_{\mathrm{s}}}$. $AA_{\rm TS}$ can be evaluated both with train or test samples and the privacy loss indicator is defined as ${\rm Privacy\ loss}=AA_{\rm TS}^{\rm test}-AA_{\rm TS}^{\rm train}$ and is expected to be strictly positive. Fig. 10 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in Yelmen et al. (2021; 2023).
@@
Figure 9: Evolution of the AATS (first row) and privacy loss (second row) indicator during the training of pretrained RBMs. The online estimation of the train and test LL is given at the third row. The overfitting is clearly detected in the HGD and MICKEY dataset and the divergence between the train and test LL coincides with a sharp rise in the privacy loss.
@@
Figure 10: Comparison of the evolution of overfitting metrics along the log-likelihood between a pretrained RBM (A) and a RBM trained from scratch using PCD (B). (C) shows the comparison between the AASyn and AATruth between the two machines. The machines were selected in order to have a comparable privacy loss (D), right before starting to overfit.
Figure 11: Comparison of log-likelihood estimation w.r.t. the exact ll on the MNIST-01 dataset.
Figure 12: Comparison of log-likelihood estimation w.r.t. the exact ll on the MNIST dataset.
Figure 13: Comparison of log-likelihood estimation w.r.t. the exact ll on the Ising dataset.
Figure 14: Comparison of log-likelihood estimation w.r.t. the exact ll on the MICKEY dataset.
Appendix I Extra figures

### v3

@@ Appendix I Overfitting and privacy loss as quality indicators
In this section, we examine the quality of the samples generated, regarding overfitting and privacy criteria which have been defined for genomic data in particular. We look at this on the models trained with PCD with and without pre-training. We focus on the human genome dataset, as shown in Fig. 1–C, to evaluate the ability of various state-of-the-art generative models to generate realistic fake genomes while minimizing privacy concerns (i.e., reducing overfitting). Recent studies (yelmen2021creating; yelmen2023deep) have thoroughly investigated this for a variety of generative models. Both studies concluded that the RBM was the most accurate method for generating high-quality and private synthetic genomes. The comparison between models relies primarily on the Nearest Neighbor Adversarial Accuracy ($AA_{\rm TS}$) and privacy loss indicators, introduced in Ref. (yale2020generation), which quantify the similarity and the level of ”privacy” of the data generated by a model w.r.t. the training set. We have $AA_{\rm TS}=\frac{1}{2}\bigl(AA_{\rm True}+AA_{\rm Synth}\bigr)$ where $AA_{\rm True}$ [resp. $AA_{\rm Synth}$] are two quantities in $[0,1]$ obtained by merging two sets of real and synthetic data of equal size $N_{s}$ and measuring respectively the frequency that a real [rep. synthetic] has a synthetic [resp. real] as nearest neighbor. If the generated samples are statistically indistinguishable from real samples, both frequencies $AA_{\rm True}$ and $AA_{\rm Synth}$ should converge to 0.5 at large $N_{\mathrm{s}}$}. $AA_{\rm TS}$ can be evaluated both with train or test samples and the privacy loss indicator is defined as ${\rm Privacy\ loss}=AA_{\rm TS}^{\rm test}-AA_{\rm TS}^{\rm train}$ and is expected to be strictly positive. Fig. 14 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in (yelmen2021creating).
@@
Figure 13: Evolution of the AATS (first row) and privacy loss (second row) indicator during the training of pretrained RBMs. The online estimation of the train and test $L$ is given at the third row. The overfitting is clearly detected in the HGD and MICKEY dataset and the divergence between the train and test $L$ coincides with a sharp rise in the privacy loss.
@@
Figure 14: Comparison of the evolution of overfitting metrics along the log-likelihood between a RBM trained from scratch using PCD (A) and a pretrained RBM (B). (C) shows the comparison between the AASyn and AATruth between the two machines. The machines were selected in order to have a comparable privacy loss (D), right before starting to overfit.
Figure 15: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the MNIST-01 dataset.
Figure 16: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the MNIST dataset.
Figure 17: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the Ising dataset.
Figure 18: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the MICKEY dataset.
Appendix J Extra figures